OpenStudy (anonymous):

MEDAL WILL BE GIVEN(: (-3/4)x + 2 = (1/2)x + 6 a. -3.2 b. 5/4 c. 16 d. 4 Also, could you explain to me how you got the answer please? Thank you in advance. I can't figure out how to do these equations with fractions.

OpenStudy (jdoe0001):

well... do you know how to add fractions?

OpenStudy (anonymous):

first you want to isolate the x. subtract 6 from both sides. then, add 3x/4 to both sides. You then are left with -4=5x/4. Multiply both sides by 4. 5x=-16. Divide both sides by 5. x=-3.2, or 3 and 1/5

OpenStudy (anonymous):

Your answer is A

OpenStudy (cj49):

(-3/4)x + 2 = (1/2)x + 6 lets look at d LHS first (-3/4)x + 2 -3x+8/4 for RHS x+12/2 -3x+8/4=x+12/2 cross multiply 2(-3x+8)=4(x+12) -6x+24=4x+48 -10x=32 x=-(32/10)=-3.2

OpenStudy (anonymous):

Thank you all so much.

OpenStudy (cj49):

Any time...